# An Adaptive Multi-Agent Bitcoin Trading System

### v1

@@ Abstract
This paper presents a Multi Agent Bitcoin Trading system that utilizes Large Language Models (LLMs) for alpha generation and portfolio management in the cryptocurrencies market. Unlike equities, cryptocurrencies exhibit extreme volatility and are heavily influenced by rapidly shifting market sentiments and regulatory announcements, making them difficult to model using static regression models or neural networks trained solely on historical data [53]. The proposed framework overcomes this by structuring LLMs into specialised agents for technical analysis, sentiment evaluation, decision-making, and performance reflection. The system improves over time through a novel verbal feedback mechanism where a Reflect agent provides daily and weekly natural-language critiques of trading decisions. These textual evaluations are then injected into future prompts, allowing the system to adjust indicator priorities, sentiment weights, and allocation logic without parameter updates or finetuning. Back-testing on Bitcoin price data from July 2024 to April 2025 shows consistent outperformance across market regimes: the Quantitative agent delivered over 30% higher returns in bullish phases and 15% overall gains versus buy-and-hold, while the sentiment-driven agent turned sideways markets from a small loss into a gain of over 100%. Adding weekly feedback further improved total performance by 31% and reduced bearish losses by 10%. The results demonstrate that verbal feedback represents a new, scalable, and low-cost method of tuning LLMs for financial goals.
@@ 1 Introduction
This paper introduces an Agentic AI system for managing a Bitcoin-only portfolio powered by an LLM : DeepSeek-r1 [31]. Conventional machine learning methods, such as regression models and neural networks, perform well in equity markets where recurring historical patterns are linked to fundamentals such as earnings, cash flows, and balance sheet strength. In contrast, Bitcoin’s price is largely driven by non-fundamental forces such as regulatory announcements, influencer statements, and speculative behavior by investors driven by fear and greed. For example, in early January 2021 Bitcoin rose from about \$29,000 to over \$40,000 in less than a week following positive commentary and Tesla’s \$1.5 billion purchase, while in May 2021 it fell nearly 30% within a week after Tesla suspended Bitcoin payments and Chinese regulators reiterated restrictions.
In such conditions, being frequently correct does not necessarily translate into making money. While predictive accuracy is intuitively linked to profitability, the correlation between the two is weak in the case of Bitcoin, which is characterized by asymmetric returns, fat tails, and high noise-to-signal ratios. Theoretically, if a model consistently predicts the correct direction of returns, profitability should follow. However, Accuracy treats all predictions equally, whereas profitability depends on the magnitude and timing of those predictions. A model that is ”often right” can still underperform if it misses major rallies or fails to exit before sharp drawdowns, while another with lower accuracy can outperform by simply capturing large moves. In such conditions, profit depends more on capital allocation, position sizing, and risk management than on raw directional correctness. Moreover, even highly accurate models may fail if execution costs, slippage, or regime shifts negate expected returns, as liquidity and volatility conditions in Bitcoin change rapidly.
@@
This promise, however, hinges on how we enable models to evolve without full retraining. Since LLMs do not inherently update their internal knowledge, the key question is how we can refresh or refine their understanding as markets change. Existing research explores methods such as Retrieval-Augmented Generation (RAG) [38], Supervised Fine-Tuning (SFT) [39], Reinforcement Learning (RL) [42], and Multi-Head Reinforcement Learning (MHRL) [41]. , but these approaches often suffer from information loss when complex human feedback is compressed into a single scalar reward signal. Several studies, including InstructGPT [42] and Constitutional AI [43], note that converting rich human feedback into a single scalar reward often omits contextual nuance and can cause models to optimize for the reward itself rather than the true intent of the feedback. To address this limitation, recent studies propose using verbal feedback: natural-language critiques provided directly to the model during inference, allowing it to retain the full semantic richness of the feedback. For instance, the Reflexion framework enables models to store and reuse their own self-critiques without any parameter updates, leading to more consistent reasoning and improved task success rates across multiple benchmarks [44]. Similarly, Self-Refine introduces an iterative “generate → critique → revise” process, where the model learns from its own feedback to progressively enhance output quality [45]. Both studies report substantial performance gains over standard LLM baselines without additional training or fine-tuning. This is especially relevant for crypto trading, where context and justification matter as much as raw predictions.
@@
To build on these capabilities, the framework structures LLMs into specialised agents for technical analysis, sentiment assessment, portfolio decision-making, and strategy evaluation. Adaptability is introduced through a novel dual verbal feedback mechanism in which the Reflect agent produces natural-language evaluations of each agent’s performance at daily and weekly intervals. The trading decisions are evaluated against benchmarks such as a static baseline buy-and-hold portfolio, risk-adjusted measures such as the Sharpe ratio, and profitability indicators such as cumulative returns, regret, and accuracy. These evaluations are then injected into the next day’s or week’s prompts. Rather than adjusting model parameters or retraining, the agents incorporate this feedback as additional context in their reasoning, allowing them to refine indicator prioritisation, sentiment weighting, and allocation choices in subsequent decisions. The system is evaluated not on classification accuracy alone but on its ability to deliver consistent profitability over time, aligning its objectives more closely with real-world investment outcomes.
@@ 2.1 Bitcoin
2.1.1 Price
Bitcoin has grown into the world’s most valuable cryptocurrency (Figure 1), breaching \$115,000 in 2025 and attaining a market capitalization of roughly \$2.3 trillion with daily trading volumes exceeding \$61 billion. Its value is sustained by a fixed supply of 21 million coins, a feature often compared to gold, and it has become widely adopted both by retail participants and large institutions. Recent milestones including the approval of spot Bitcoin ETFs and its use as a corporate treasury asset have cemented its role as the benchmark of the digital asset market [47]. Over the last decade, Bitcoin has outperformed traditional assets by an extraordinary margin, growing roughly 26,900% versus $\sim$193% for the S&P 500 and $\sim$126% for gold, with a 10-year CAGR for Bitcoin of 124% compared to 10–13% for equities and gold.
@@ 2.1.1 Price
Figure 2: Historical Rolling Annualized Volatility (Standard Deviation) Comparison 2015-2025 for Gold, S&P 500, and Bitcoin
This stark contrast in volatility illustrates that Bitcoin’s price movements are far more unpredictable and prone to sharp swings, making it a substantially higher risk asset relative to the traditional safe haven status of Gold and the diversified equity exposure of the S&P 500. From an investor’s standpoint, Bitcoin rewards those with a high risk tolerance and a long horizon, while Gold is best suited for capital preservation and cushioning drawdowns. The S&P 500 remains the core growth anchor for most portfolios. When combined, these three assets can complement one another Bitcoin driving upside potential, Gold adding stability, and the S&P 500 delivering steady growth resulting in a diversified mix that manages risk while seeking attractive returns. Ultimately, the right balance depends on an investor’s risk appetite and time horizon.
@@ 2.1.3 Price Factors
Bitcoin’s price is influenced by a combination of sentiment, technical, and on-chain factors that together capture both market psychology and structural dynamics [55]. Sentiment plays a decisive role in short-term movements, as shifts in public mood driven by regulation, macroeconomic events, or social media discussions often lead to rapid rallies or declines. Technical indicators derived from price and volume, such as moving averages, momentum oscillators, and volatility bands, provide a systematic way to identify trends, turning points, and risk conditions within this volatility. Alongside these, on-chain data offers a unique perspective by measuring the actual use of the network, with metrics such as transaction volumes and active addresses serving as signals of adoption and liquidity. Each of these dimensions is important: sentiment explains sudden market swings, technical analysis offers structured decision rules, and on-chain metrics capture the underlying health of the system. Together, they form a comprehensive framework for understanding and anticipating Bitcoin’s price behaviour.
@@ 2.2 Large Language Models in Finance
The use of LLMs in finance is not without risk. Three persistent limitations constrain their deployment in trading contexts. First, they are prone to hallucination: generating plausible but false information. In one benchmark, a finance-focused LLM produced fabricated content in 41% of probing test cases, raising concerns about reliability [49]. Second, their knowledge is temporally bounded by training cutoffs, making them poorly suited to tasks requiring real-time awareness unless supplemented by retrieval or live feeds. Third, they lack interpretability. With billions of parameters, it is often unclear why a particular recommendation is produced, creating difficulties for accountability in financial decision-making. Together, these weaknesses necessitate the design of auxiliary mechanisms to ensure robustness, adaptability, and transparency.
@@ 3 System Design and Methodology
The proposed system is designed not only to evaluate profitability in a portfolio across different market regimes but also to identify the contribution of individual market factors such as technical indicators and sentiment signals in generating profitable trades.
@@ 3.1 Quantitative Agent (Quants)
It relies on structured, market-derived data to ground its predictions. This includes OHLCV data (open, high, low, close, and volume) drawn from exchange feeds, as well as a range of technical indicators computed from this data, such as moving averages, MACD, RSI, Bollinger Bands, VWAP, and ADX. To complement these price-based measures, the agent also integrates on-chain metrics sourced directly from the Bitcoin blockchain, including transaction counts, active addresses, and transfer volumes. These inputs ensure that the Quants Agent captures both price mechanics and underlying network activity, enabling it to generate daily market state predictions and suggested portfolio allocations. By combining exchange-based and blockchain-native data, the agent is positioned to balance short-term technical signals with deeper structural insights into adoption and liquidity.
@@ 3.2 Sentiment Agent (Signals)
The Signals Agent is responsible for assessing investor sentiment related to Bitcoin. It draws on three primary data sources that capture market psychology and external influences on Bitcoin. First, news data from trusted financial outlets (e.g., CNBC, Yahoo Finance, Forbes, Business Insider, Seeking Alpha, and Fortune) is aggregated via the GNews API, ensuring coverage of regulatory announcements, macroeconomic shifts, and key narratives shaping sentiment. Second, the Fear and Greed Index, obtained from Alternative.me, provides a composite daily score ranging from extreme fear to extreme greed, reflecting crowd mood based on volatility, trading volume, social media, and surveys. Finally, social media sentiment scores are collected through the Senticrypt API, which processes Twitter discussions to quantify retail investor mood in real time. Together, these inputs allow the agent to evaluate both structured and unstructured sentiment signals, ensuring that its predictions incorporate news-driven shocks, aggregated market psychology, and the daily pulse of social platforms. The classification and reasoning are passed to the Decision Agent, while the portfolio is forwarded to the Reflect Agent for evaluation.
@@ 3.4.1 Reflect Agent (Daily Feedback)
Prompt tuning in this system was particularly challenging because feedback generated by the Reflect Agent often strayed beyond the intended scope. Early experiments showed that the Signals Agent, which is restricted to sentiment and news data, was incorrectly instructed to incorporate technical indicators – data it did not have access to. To correct this, prompts had to be explicitly constrained so that feedback remained independent and tied only to the inputs available to each agent. A similar issue arose when feedback provided direct portfolio advice, such as suggesting the Decision Agent increase its Bitcoin allocation by 10%. Allowing such instructions would undermine the integrity of the reasoning process, as decisions must emerge from data-driven analysis rather than arbitrary adjustments. To address this, prompts were carefully designed to ensure that feedback focused only on critiquing reasoning, identifying errors, and suggesting improvements in analysis, without dictating specific allocations. These constraints made the system significantly harder to scale, since adding more agents or expanding their roles requires increasingly precise prompt design to prevent cross-contamination of inputs and preserve the validity of the reasoning process.
@@ 3.4.2 Long Term Reflect Agent (Weekly Feedback)
No autonomous strategy can sustain profitability in the Bitcoin market without some form of oversight. This agent is therefore designed to supply a systematic analogue of that oversight: a weekly evaluation mechanism that mirrors the role of a human supervisor by reviewing agent performance over time, identifying deviations from expected behaviour, and suggesting appropriate corrective actions.
However, daily oversight alone is insufficient. The Reflect Agent provides day-by-day feedback, but single-day evaluations cannot capture slower-moving trends or recurring issues, and the system’s zero-shot prompting means there is no built-in memory across days. Although agents can react to the previous day’s notes, they do not retain awareness of mistakes made earlier in the week or month. To address this, the Long-Term Feedback Agent condenses a full trading week into a single summary that is persisted throughout the following week, effectively compensating for the stateless nature of the language models and supplying a structured form of memory. This weekly context also maintains a broader view of market conditions, guiding and strengthening the agents’ reasoning over time.

### v2

@@ Abstract
This paper presents a Multi Agent Bitcoin Trading system that utilizes Large Language Models (LLMs) for alpha generation and portfolio management in the cryptocurrencies market. Unlike equities, cryptocurrencies exhibit extreme volatility and are heavily influenced by rapidly shifting market sentiments and regulatory announcements, making them difficult to model using static regression models or neural networks trained solely on historical data. The proposed framework overcomes this by structuring LLMs into specialised agents for technical analysis, sentiment evaluation, decision-making, and performance reflection. The agents improve over time via a novel verbal feedback mechanism where a Reflect agent provides daily and weekly natural-language critiques of trading decisions. These textual evaluations are then injected into future prompts of the agents, allowing them to adjust allocation logic without weight updates or finetuning. Back-testing on Bitcoin price data from July 2024 to April 2025 shows consistent outperformance across market regimes: the Quantitative agent delivered over 30% higher returns in bullish phases and 15% overall gains versus buy-and-hold, while the sentiment-driven agent turned sideways markets from a small loss into a gain of over 100%. Adding weekly feedback further improved total performance by 31% and reduced bearish losses by 10%. The results demonstrate that verbal feedback represents a new, scalable, and low-cost approach of tuning LLMs for financial goals.
@@ 1 Introduction
This paper introduces an Agentic AI system for managing a Bitcoin-only portfolio powered by DeepSeek-r1 [5]. Conventional machine learning methods, such as regression models and neural networks, perform well in equity markets where recurring historical patterns are linked to fundamentals such as earnings, cash flows, and balance sheet strength  [23]. In contrast, Bitcoin’s price is largely driven by non-fundamental forces such as regulatory announcements, influencer statements, and speculative behavior by investors driven by fear and greed. For example, in early January 2021 Bitcoin rose from about \$29,000 to over \$40,000 in less than a week following positive commentary and Tesla’s \$1.5 billion purchase, while in May 2021 it fell nearly 30% within a week after Tesla suspended Bitcoin payments and Chinese regulators reiterated restrictions. Therefore, Bitcoin is characterized by asymmetric returns, fat tails, and high noise-to-signal ratios.
In such conditions, being frequently correct does not necessarily translate into making money. Theoretically, if a model consistently predicts the correct direction of returns, profitability should follow. However, Accuracy treats all predictions equally, whereas profitability depends on the magnitude and timing of those predictions. A model that is ”often right” can still underperform if it misses major rallies or fails to exit before sharp drawdowns, while another with lower accuracy can outperform by simply capturing large moves. In such conditions, profit depends more on capital allocation, position sizing, and risk management than on raw directional correctness. Moreover, even highly accurate models may fail if execution costs, slippage, or regime shifts negate expected returns.
@@
This promise, however, hinges on how we enable LLMs to evolve without full retraining. Since LLMs do not inherently update their internal knowledge, the key question is how we can refresh or refine their understanding as markets change. Existing research explores methods such as Retrieval-Augmented Generation (RAG) [8], Supervised Fine-Tuning (SFT) [9], Reinforcement Learning (RL) [12], and Multi-Head Reinforcement Learning (MHRL) [11]. , but these approaches often suffer from information loss when complex human feedback is compressed into a single scalar reward signal. Several studies, including InstructGPT [12] and Constitutional AI [13], note that converting rich human feedback into a single scalar reward often omits contextual nuance and can cause models to optimize for the reward itself rather than the true intent of the feedback. To address this limitation, recent studies propose using verbal feedback: natural-language critiques provided directly to the model during inference, allowing it to retain the full semantic richness of the feedback. For instance, the Reflexion framework enables models to store and reuse their own self-critiques without any parameter updates, leading to more consistent reasoning and improved task success rates across multiple benchmarks [14]. Similarly, Self-Refine introduces an iterative “generate → critique → revise” process, where the model learns from its own feedback to progressively enhance output quality [15]. Both studies report substantial performance gains over standard LLM baselines without additional training or fine-tuning. This is especially relevant for crypto trading, where context and justification matter as much as raw predictions.
Now that this system incorporates natural-language critiques as its optimisation method, the next step is to examine how these capabilities translate into actual trading performance. Improving a trading strategy depends on identifying which input factors matter most for Bitcoin. Is it technical indicators, sentiment measures, or broader market variables? The existing literature remains inconclusive on this point, with studies reaching mixed conclusions about which features provide consistently useful signals.
@@
To build on these capabilities, the framework structures LLMs into specialised agents for technical analysis, sentiment assessment, portfolio decision-making, and strategy evaluation. Adaptability is introduced through a novel dual verbal feedback mechanism in which the Reflect agent produces natural-language evaluations of each agent’s performance at daily and weekly intervals. The trading decisions are evaluated against benchmarks such as a static baseline buy-and-hold portfolio, risk-adjusted measures such as the Sharpe ratio, and profitability indicators such as cumulative returns, regret, and accuracy. These evaluations are then injected into the next day’s or week’s prompts. Rather than adjusting model parameters or retraining, the agents incorporate this feedback as additional context in their reasoning, allowing them to refine indicator prioritisation, sentiment weighting, and allocation choices in subsequent decisions. Each agent manages its own portfolio based on distinct inputs, such as quantitative indicators or sentiment data, allowing a direct comparison of which factors lead to stronger trading performance. The system is evaluated not on classification accuracy alone but on its ability to deliver consistent profitability over time, aligning its objectives more closely with real-world investment outcomes.
@@ 2.1 Bitcoin
Bitcoin has grown into the world’s most valuable cryptocurrency (Figure 2), breaching \$115,000 in 2025 and attaining a market capitalization of roughly \$2.3 trillion with daily trading volumes exceeding \$61 billion. Its value is sustained by a fixed supply of 21 million coins, a feature often compared to gold, and it has become widely adopted both by retail participants and large institutions. Recent milestones including the approval of spot Bitcoin ETFs and its use as a corporate treasury asset have cemented its role as the benchmark of the digital asset market [17].
2.1.1 Price
@@
Figure 2: Historical Price Volatility (2020–2025) for Gold, S&P 500, and Bitcoin
@@
Over the last decade, Bitcoin has outperformed traditional assets by an extraordinary margin, growing roughly 26,900% versus $\sim$193% for the S&P 500 and $\sim$126% for gold, with a 10-year CAGR for Bitcoin of 124% compared to 10–13% for equities and gold as seen in Table  1.
But the CAGR doesn’t explain the stark contrast in volatility as shown in Figure  2. Bitcoin’s price movements are far more unpredictable and prone to sharp swings, making it a substantially higher risk asset relative to the traditional safe haven status of Gold and the diversified equity exposure of the S&P 500. From an investor’s standpoint, Bitcoin rewards those with a high risk tolerance and a long horizon, while Gold is best suited for capital preservation and cushioning drawdowns. The S&P 500 remains the core growth anchor for most portfolios. When combined, these three assets can complement one another Bitcoin driving upside potential, Gold adding stability, and the S&P 500 delivering steady growth resulting in a diversified mix that manages risk while seeking attractive returns. Ultimately, the right balance depends on an investor’s risk appetite and time horizon.
@@ 2.1.3 Price Factors
Bitcoin’s price is mainly shaped by how limited its supply is, how investors behave, and how active the network is. The total number of bitcoins is capped at 21 million, and the reward for mining new coins is cut in half every few years. This built-in scarcity means that when more people want to buy or use bitcoin, its price tends to rise quickly. According to Investopedia, bitcoin’s value is driven by its supply, demand, competition from other cryptocurrencies, and investor sentiment.
In the short term, prices move mostly because of changes in mood and expectations. News about regulation, economic trends, or social media discussions can quickly shift how investors feel, leading to sharp gains or drops. These swings often feed on themselves : optimism pushes prices higher, while fear causes sudden sell-offs (Investopedia).
Recent research also shows that technical indicators remain useful for understanding short-term movements in bitcoin. Empirical studies find that price and volume-based indicators often contain statistically significant predictive information in cryptocurrency markets. [26] shows that incorporating standard technical indicators into deep-learning models improves the accuracy of crypto price forecasts relative to models using raw prices alone. Similarly, [27] finds that classical indicators, such as moving averages and stochastic oscillators, enhance prediction performance across multiple cryptocurrencies when evaluated with machine-learning methods. More recently, [28] demonstrates that a range of technical trading rules can outperform a simple buy-and-hold benchmark under certain market conditions, indicating that technical signals still capture meaningful short-term structure in crypto price dynamics.
Network data also plays an important role. Information such as how many transactions are happening, how many users are active, and how miners behave shows how much the system is being used in practice. Combining this with market and sentiment data gives a clearer picture of what drives bitcoin’s price and whether the movements are based on real growth or short-term speculation. Together, these findings suggest that well-designed technical indicators, market sentiment, and on chain variables continue to provide practical insight into near-term bitcoin price behaviour.
@@ 2.2 Large Language Models in Finance
However, The use of LLMs in finance is not without risk. Three persistent limitations constrain their deployment in trading contexts. First, they are prone to hallucination: generating plausible but false information. In one benchmark, a finance-focused LLM produced fabricated content in 41% of probing test cases, raising concerns about reliability [19]. Second, their knowledge is temporally bounded by training cutoffs, making them poorly suited to tasks requiring real-time awareness unless supplemented by retrieval or live feeds. Third, they lack interpretability. With billions of parameters, it is often unclear why a particular recommendation is produced, creating difficulties for accountability in financial decision-making. Together, these weaknesses necessitate the design of auxiliary mechanisms to ensure robustness, adaptability, and transparency.
@@
A balanced view therefore recognises both the capabilities and constraints of LLMs in financial settings. Their strength lies in processing heterogeneous information and adapting to new tasks without retraining, yet these advantages come with risks such as hallucination, improper instruction following, and lack of interpretability. Effective deployment requires careful prompt design, structured interaction protocols, and mechanisms that constrain model behaviour.
Figure 5: System Architecture
@@ 3 System Design and Methodology
Building on the findings of the background research, the proposed system is designed not only to evaluate profitability in a portfolio across different market regimes but also to identify the contribution of individual market factors in generating profitable trades. The system architecture can be seen in figure  5. The system prompts for each agent can be found in the appendix.
@@ 3.1 Quantitative Agent (Quants)
It relies on structured, market-derived data to ground its predictions. This includes OHLCV data (open, high, low, close, and volume) drawn from exchange feeds  [4], as well as a range of technical indicators computed from this data using the Python TA module, such as moving averages, MACD, RSI, Bollinger Bands, VWAP, and ADX. To complement these price-based measures, the agent also integrates on-chain metrics sourced directly from the Bitcoin blockchain  [7], including transaction counts, active addresses, and transfer volumes. These inputs ensure that the Quants Agent captures both price mechanics and underlying network activity, enabling it to generate daily market state predictions and suggested portfolio allocations. By combining exchange-based and blockchain-native data, the agent is positioned to balance short-term technical signals with deeper structural insights into adoption and liquidity.
@@ 3.2 Sentiment Agent (Signals)
The Signals Agent is responsible for assessing investor sentiment related to Bitcoin. It draws on three primary data sources that capture market psychology and external influences on Bitcoin. First, news data from trusted financial outlets (e.g., CNBC, Yahoo Finance, Forbes, Business Insider, Seeking Alpha, and Fortune) is aggregated via the GNews API  [3], ensuring coverage of regulatory announcements, macroeconomic shifts, and key narratives shaping sentiment. Second, the Fear and Greed Index, obtained from Alternative.me, provides a composite daily score ranging from extreme fear to extreme greed, reflecting crowd mood based on volatility, trading volume, social media, and surveys  [1]. Finally, social media sentiment scores are collected through the Senticrypt API, which processes Twitter discussions to quantify retail investor mood in real time  [2].Together, these inputs allow the agent to evaluate both structured and unstructured sentiment signals, ensuring that its predictions incorporate news-driven shocks, aggregated market psychology, and the daily pulse of social platforms. The classification and reasoning are passed to the Decision Agent, while the portfolio is forwarded to the Reflect Agent for evaluation.
@@ 3.4.2 Long Term Reflect Agent (Weekly Feedback)
No autonomous strategy can sustain profitability in the Bitcoin market without some form of oversight. While the Reflect Agent provides day-by-day evaluations, single-day feedback cannot capture slower behavioural drift, recurring mistakes, or emerging patterns in agent behaviour. Because the system operates through zero-shot prompting, agents do not retain memory of issues identified earlier in the week. A mechanism is therefore required to supply a broader temporal perspective : one that reviews behaviour over multiple days, identifies persistent weaknesses, and maintains continuity that the underlying language models cannot provide on their own. The weekly evaluation module fulfils this role by functioning as a supervisory layer that consolidates weekly performance and anchors agent behaviour within a longer horizon.
The important detail is that this weekly oversight is not powered by an LLM. Instead, the system uses a deterministic rule-based evaluator that produces agent-specific recommendations. Each agent is assessed independently based on its regret, Sharpe ratio, and realised returns for the week. For each metric, the evaluator first determines whether performance is positive or negative and then assigns a severity level (mild, strong, exceptional). The corresponding suggestion is then selected from the predefined rule set. For example, strong positive regret for the quant agent returns advice to maintain its current reasoning while adapting to shifting conditions, while an exceptional negative Sharpe ratio for the signals agent triggers guidance to re-evaluate its source list and reduce volatility
The weekly recommendation that results from this evaluation is then carried forward into the following week and included in every daily prompt. By persisting the prior week’s summary in this way, the system provides the agents with a consistent supervisory signal across time, reinforcing long-horizon insights and ensuring that previous weaknesses are not forgotten as market conditions evolve. The recommendations can be found in the appendix.
3.4.3 Prompt Engineering
Prompt engineering is central to this system because each agent must receive clear, contextual, and precise instructions, with an explicit role definition. These elements ensure that agents interpret constraints consistently and generate reasoning traces that downstream components can evaluate. Recent surveys on prompt engineering show that structured instructions, repeated priorities, and well-defined roles improve controllability in complex workflows [29, 30]. These principles motivate the formulation of each system prompt. Essential rules are restated and emphasised using terms such as highly important to reduce forgetting across long reasoning sequences. The system prompts for all the agents are provided in the appendix.
Additional constraints were required for the daily feedback agent because it’s output drives improvements in the trading performances of the agents. Early experiments showed that the Reflect Agent occasionally produced feedback outside its scope, despite explicit role definition. For example, the Signals Agent, which only receives sentiment and news inputs, was instructed to incorporate technical indicators it could not access. Similar issues occurred when feedback included explicit portfolio adjustments for the Decision Agent, such as recommending a fixed allocation change. Allowing such directives would break the reasoning pathway because allocation decisions must follow from the observed data rather than external prescriptions. The prompts were therefore refined so that feedback evaluates reasoning quality, identifies errors, and proposes analytical improvements without dictating actions or referencing unavailable inputs. These constraints introduce a scaling challenge, as adding new agents or expanding existing roles requires precise prompt structure to prevent information leakage and maintain interpretability of the full decision process.